Five years ago, Liz was twice as old as her younger sister. Five years from now, Liz will only be 1 1/3 times as old as her sister. How old is each now?
I'm really confused could someone help me please?

Answer by stanbon(75887)

Five years ago, Liz was twice as old as her younger sister. Five years from now, Liz will only be 1 1/3 times as old as her sister. How old is each now?
--------------------
Five years ago DATA:
Let sister be "x" yrs. old ; Liz was "2x" yrs. old
--------------
Ages NOW DATA:
Sister is x+5 yrs. old ; Liz is "2x+5" yrs. old.
----------------
Five years from NOW DATA
Sis will be x+10 yrs old ; Liz will be 2x+10 yrs old
-----------------
EQUATION:
Five years from now, Liz will only be 1 1/3 times as old as her sister.
2x+10 = (4/3)(x+10)
2x-(4/3)x = (40/3)-10
(2/3)x = 10/3
x = 5 (Sis age NOW)
2x = 10 (Liz age NOW)
==============
Cheers,
Stan H.